# Takahiro Nishinakaa,b,c and Yutaka Yoshida d,e

### v1

@@ Appendix C Gauge and R-charge for dressed monopole operator
Here we summarize the formulas, for example see [49], for the gauge charge and $R$-charges of monopole operators, dressed by chiral multiplets. Again let us consider an abelian $\mathcal{N}=2$ CS matter theory with the gauge group $U(1)^{N}$ coupled to $L$ chiral multiplets $\{\phi_{i}\}_{i=1}^{L}$ with the R-charge $r_{i}$. We denote the gauge CS level by $K_{ab}$ with $a,b=1,\cdots,N$, and denote the gauge-R-symmetry mixed CS level by $K_{aR}$. We also denote the gauge charge of the chiral multiplets by $Q_{ai}$ with $a=1,\cdots,N$ and $i=1,\cdots,L$.
We consider a BPS dressed monopole operator $\mathcal{O}_{(n,m)}=(\prod_{i=1}^{L}\phi^{n_{i}}_{i})V_{m}$ with magnetic charge $m=(m_{1},\cdots,m_{N})\in\mathbb{Z}^{N}$. Here $nn_{i}$ has to satisfy $n_{i}(m\cdot Q_{i})=0$ for $i=1,\cdots,L$, and $m\cdot Q_{i}:=\sum_{a=1}^{N}m_{a}Q_{ai}$.
@@
$Q_{a}[\mathcal{O}_{(n,m)}]=\sum_{b=1}^{N}K_{ab}m_{b}+\sum_{i=1}^{L}n_{i}Q_{ai}-\sum_{i=1}^{L}\frac{1}{2}Q_{ai}(|m\cdot Q_{i}|+{m}\cdot Q_{i}),$ (C.1)

### v2

@@ Appendix C Gauge and R-charge for dressed monopole operator
Here we summarize the formulas, for example see [50], for the gauge charge and $R$-charges of monopole operators, dressed by chiral multiplets. Again let us consider an abelian $\mathcal{N}=2$ CS matter theory with the gauge group $U(1)^{N}$ coupled to $L$ chiral multiplets $\{\phi_{i}\}_{i=1}^{L}$ with the R-charge $r_{i}$. We denote the gauge CS level by $K_{ab}$ with $a,b=1,\cdots,N$, and denote the gauge-R-symmetry mixed CS level by $K_{aR}$. We also denote the gauge charge of the chiral multiplets by $Q_{a,i}$ with $a=1,\cdots,N$ and $i=1,\cdots,L$.
We consider a BPS dressed monopole operator $\mathcal{O}_{(n,m)}=(\prod_{i=1}^{L}\phi^{n_{i}}_{i})V_{m}$ with magnetic charge $m=(m_{1},\cdots,m_{N})\in\mathbb{Z}^{N}$. Here $n_{i}$ has to satisfy $n_{i}(m\cdot Q_{i})=0$ for $i=1,\cdots,L$, and $m\cdot Q_{i}=\sum_{a=1}^{N}m_{a}Q_{a,i}$.
@@
$Q_{a}[\mathcal{O}_{(n,m)}]=\sum_{b=1}^{N}K_{ab}m_{b}+\sum_{i=1}^{L}n_{i}Q_{a,i}-\sum_{i=1}^{L}\frac{1}{2}Q_{a,i}(|m\cdot Q_{i}|+{m}\cdot Q_{i}),$ (C.1)